OpenStudy (abmon98):

From the top of a vertical tower 248 m high a stone is projected vertically downwards with a speed of 8ms^-1. After x seconds, another stone is projected vertically upwards from the level of the base of the tower with a speed of 25ms^-1.Given that the two stones are at the same height moving in the same direction, 6 second after the projection of the first stone, calculate a) The value of x b) The velocity of the second stone at this instant

OpenStudy (anonymous):

s1 = -8 t - (1/2) (9.8 m/s^2) t^2 s2 = 25 (t-x) -(1/2)(9.8 m/s^2)(t-x)^2 s1 = s2 at t=6 substitute and solve for x. v2 = 25 - (9.8 m/s^2)(t-x)

OpenStudy (abmon98):

why is it -8

OpenStudy (anonymous):

correction: x = x0 + v0 t + (1/2) a t^2

OpenStudy (nincompoop):

it's going down

OpenStudy (abmon98):

so the distance that the first stone covered is 228m

OpenStudy (abmon98):

so 248-228=20

OpenStudy (abmon98):

so 20=25(6-x)-1/2(10)(6-x)^2=-5x^2+35x-50=0

OpenStudy (abmon98):

x= either 2 or 5 so 6-2=4 seconds

OpenStudy (abmon98):

use v^2=u^2+2as so (25*25+2(-10)(20)^1/2

OpenStudy (abmon98):

i checked out the answers in my mechanics book and it appears to be 2 and 14.2

OpenStudy (abmon98):

when i replaced the t with 6 in the last part of your working i have found out the answer to be 458

OpenStudy (abmon98):

anyways thank you for your help

OpenStudy (abmon98):

and i know that it decelerates, what iam saying here is that when i substituted t by 6 the answer is not still the same as in the markscheme of my book

OpenStudy (nincompoop):

corrected position function s(t) = -4.5t^2 - 8t + 248

OpenStudy (nincompoop):

@ganeshie8 @agent0smith

OpenStudy (nincompoop):

evaluate at s(6) and this should reveal to you where the stone is

OpenStudy (nincompoop):

@ganeshie8 how would you set up the second stone with the time delay? I was just going to do a straight forward s(t) = -4.5t^2 + 25t + 0 then evaluate at s(6) also then figure out where the position is supposed to be obviously this will be different because it has a greater initial velocity and it is decelerating as it goes up. after evaluating at s(6), solve for t at the distance where the two stones meet add (or subtract) the t to 6

ganeshie8 (ganeshie8):

\(s(t_1) = -4.5t_1^2 - 8t_1 + 248\) \(s(t_2) = -4.5t_2^2 + 25t_2 + 0\)

ganeshie8 (ganeshie8):

and since second stone is fired x time after first stone, \(t_2 = t_1 + x\)

ganeshie8 (ganeshie8):

ur method also works... but in the background u r using t2 is lagging i think...

OpenStudy (nincompoop):

I didn't know how to squeeze the delay into the second function I was thinking that the second function be evaluated how long it will take for it to reach the same position as the first stone's s(6) position 6-t = x

ganeshie8 (ganeshie8):

i made a typo earlier. corrected below : \(t_2 = t_1-x\) so that t_2 becomes 0 when t_1 = x, implying that second stone starts x seconds after the first stone